A liquid has a volume of 40.5 mL and a mass of 44.6 g . What is the specific gravity of the liquid?
1
Expert's answer
2020-12-06T10:32:43-0500

density of liquid = (mass / volume)

density of liquid = (44.6/40.5) g/mL = 1.101g/mL

density of water = 1g/mL

specific gravity = "\frac{density\ of\ liquid}{density\ of\ water}"

specific gravity = 1.101